# Primordial Black Hole Interpretation in Subsolar Mass Gravitational Wave Candidate SSM200308

### v1

@@ Abstract
In the recent second part of the third observation run by the LIGO-Virgo-KAGRA collaboration, a candidate with sub-solar mass components was reported, which we labelled as SSM200308. This study investigates the premise that primordial black holes (PBHs), arising from Gaussian perturbation collapses, could explain SSM200308. Through Bayesian analysis, we obtain the primordial curvature power spectrum that leads to the merger rate of PBHs aligning with observational data as long as they constitute $3.54^{+7.66}_{-2.82}\times 10^{-3}$ of the dark matter. However, while the gravitational wave (GW) background from binary PBH mergers is within current observational limits, the scalar-induced GWs associated with PBH formation exceed the constraints imposed by pulsar timing arrays, challenging the Gaussian perturbation collapse PBH model as the source of SSM200308.
@@ I Introduction
Since the first detection of a gravitational wave (GW) event by LIGO Abbott et al. (2016a), which marked the dawn of GW astronomy, the LIGO-Virgo-KAGRA collaboration has reported nearly a hundred GW events over the subsequent years.Abbott et al. (2016b, 2018a, 2021, 2024, 2023a). Prior investigations into compact binary coalescences of sub-solar mass (SSM) objects within the collected data yielded no conclusive evidence of such events Abbott et al. (2018b, 2019); Nitz and Wang (2021); Abbott et al. (2022, 2023b); Nitz and Wang (2022). However, it was not until recently, in the second part of the third observing run Abbott et al. (2023a), that three SSM candidates were reported Abbott et al. (2023b) and they are also analyzed in Morras et al. (2023); Prunier et al. (2023). These three candidates are not claimed as GW events, owing to their relatively large false alarm rates and relatively low signal-to-noise ratio (SNR). Notably, one of the candidates, SSM200308, exhibited a false alarm rate of approximately $0.2$ per year and an $\mathrm{SNR}=8.90$, positioning it on the cusp of claiming a GW detection. With ongoing enhancements in LIGO-Virgo-KAGRA’s sensitivity and the accumulation of observational data, the investigation into the nature and origins of SSM candidates is emerging as a promising and compelling field of study. Recent research has proposed primordial black holes (PBH) as a potential source of SSM candidates Yamamoto et al. (2023); Carr et al. (2024); Crescimbeni et al. (2024), suggesting a novel avenue for understanding these enigmatic phenomena.
PBHs are formed in the very early Universe, originating from the collapse of overmassive regions Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974); Carr (1975). The overabundance of mass in these regions is generated by large curvature perturbations that are enhanced on scales (e.g., Cai et al. (2018); Cotner and Kusenko (2017); Espinosa et al. (2018); Palma et al. (2020); Pi and Sasaki (2023); Meng et al. (2023)) much smaller than those observed in the cosmic microwave background Aghanim et al. (2020). PBHs are not only promising dark matter (DM) candidates but if a few-thousandth of DM consists of PBHs, they can also explain the GW events observed by the LIGO-Virgo-KAGRA collaboration Sasaki et al. (2016); Chen and Huang (2018); Raidal et al. (2019); De Luca et al. (2020); Hall et al. (2020); Bhagwat et al. (2021); Hütsi et al. (2020); Wong et al. (2021); De Luca et al. (2021); Bavera et al. (2022); Franciolini et al. (2021); Chen et al. (2022); Chen and Hall (2024).
@@ II Bayesian analysis of SSM200308
$\mathcal{R}_{12}(t\mid\vec{\theta})\approx 3.9\cdot 10^{6}\times\left({\frac{t}{t_{0}}}\right)^{-\frac{34}{37}}f_{\mathrm{PBH}}^{2}(0.7f_{\mathrm{PBH}}^{2}+\sigma_{\mathrm{eq}}^{2})^{-{21\over 74}}$ (1)
×$\times\min\left(\frac{P(m_{1}|\vec{\theta})}{m_{1}},\frac{P(m_{2}|\vec{\theta})}{m_{2}}\right)\left({P(m_{1}|\vec{\theta})\over m_{1}}+{P(m_{2}|\vec{\theta})\over m_{2}}\right)$
×$\times(m_{1}m_{2})^{{3\over 37}}(m_{1}+m_{2})^{36\over 37},$
where $t_{0}$ is the age of the Universe, $f_{\mathrm{PBH}}\equiv\Omega_{\mathrm{PBH}}/\Omega_{\mathrm{CDM}}$ represents the energy fraction of CDM in the form of PBHs and $\sigma_{\mathrm{eq}}\approx 0.005$ Ali-Haïmoud et al. (2017); Chen and Huang (2018) stands for the variance of density perturbations of the rest CDM at matter-radiation equality. The component masses in Eq. (1) are given in the units of $M_{\odot}$ and $\vec{\theta}$ represents a population of parameters in the PBH mass function. The mass function is normalized as $\int P(m)\mathrm{d}m=1$.
@@
where $\mathcal{C}(r)=\mathcal{C}_{g}-\frac{1}{4\Phi}\mathcal{C}_{g}^{2}$ and $\mathcal{C}_{g}=-2\Phi r\zeta^{\prime}(r)$ is linearly related to the comoving curvature perturbation. Here, $\Phi\equiv 3(1+w)/(5+3w)$ and $w$ is the equation of state. Note that during the QCD phase transition, due to the softening of the equation of state Musco and Miller (2013); Saikawa and Shirai (2018), the threshold value of forming PBHs will slightly decrease Borsanyi et al. (2016); Saikawa and Shirai (2018); Franciolini et al. (2022); Musco et al. (2023) since there is less pressure to resist the collapse overdensities. $r_{m}$ in Eq. (2) represents the distance between the location where $\mathcal{C}_{g}$ gains its maximum from and the origin of the peak in curvature perturbation, $\zeta(0)$. The mass of PBH over the horizon mass is given by $m/M_{H}=\kappa\left(\mathcal{C}(r_{m})-\mathcal{C}_{c}\right)^{\gamma}$ Choptuik (1993); Evans and Coleman (1994); Niemeyer and Jedamzik (1998) with $\kappa\simeq 3.3$ and $\gamma\simeq 0.36$ Koike et al. (1995). The second-order derivative of the compact function is given by
@@
with $\vec{\lambda}\equiv\{m_{1},m_{2}\}$, $p(\vec{\lambda}|data)$ is the posterior of SSM200308 obtained in Prunier et al. (2023) and $\beta(\vec{\theta})\equiv\int\mathrm{d}\vec{\lambda}\ VT(\vec{\lambda})\ \mathcal{R}_{12}(\vec{\lambda}|\vec{\theta})$ where $VT(\vec{\lambda})$ is the spacetime sensitivity volume of LIGO-Virgo-KAGRA detectors Abbott et al. (2023b).
@@
We report our analysis in Fig. 1. The median value of the parameters and their $95\%$ equal-tailed credible intervals are $A=0.24^{+0.42}_{-0.11}$, $k_{*}=4.53^{+1.32}_{-1.06}\times 10^{6}\mathrm{Mpc}^{-1}$ and $\sigma=0.51^{+0.40}_{-0.24}$. Moreover, fraction of PBHs in DM is evaluated to be $f_{\mathrm{PBH}}=3.54^{+7.66}_{-2.82}\times 10^{-3}$ and the local merger rate is given by $\mathcal{R}=3836^{+30906}_{-3670}\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$. The sensitivity volume values from the O3 run in the SSM search for SSM200308 is $\left\langle VT\right\rangle\simeq 3\times 10^{-4}\mathrm{Gpc}^{3}\mathrm{yr}$ Abbott et al. (2023b). Therefore, the expected event numbers are $N=\mathcal{R}\left\langle VT\right\rangle=1.05^{+8.50}_{-1.01}$ which is consistent with Phukon et al. (2021); Abbott et al. (2023b).
Furthermore, we demonstrate the PBH mass function in Fig. 2, together with the median mass of SSM200803. We also show the constraints on $f_{\mathrm{PBH}}$ by EROS/MACHO microlensing Tisserand et al. (2007), OGLE microlensing Niikura et al. (2019), SGWB from binary PBHs Wang et al. (2018); Chen and Huang (2019), dynamical heating of ultra-faint dwarf galaxies (UFD) Brandt (2016) and accretion constraints of CMB Ali-Haimoud and Kamionkowski (2017); Aloni et al. (2017); Horowitz (2016); Chen et al. (2016); Poulin et al. (2017). It can be seen that the reconstructed mass function (or the abundance that $f_{\mathrm{PBH}}=3.54^{+7.66}_{-2.82}\times 10^{-3}$) is consistent with current observational constraints.
Figure 2: The posterior distribution of the PBH mass function. The blue bands represent the region at $95\%$ confidence level. The gray dashed lines are the median mass of SSM200803. We also present the current constraints on $f_{\mathrm{PBH}}$.
@@ III.2 SGWB form binary PBH coalescences
Here $H_{0}$ is the Hubble value by today, $\Delta=\Omega_{r}(1+z)^{4}+\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}$, $\Omega_{m}$ and $\Omega_{\Lambda}$ are density parameters for matter and dark energy respectively Aghanim et al. (2020). The energy spectrum of a single PBH coalescence event in the source frame, ${dE_{s}}/{df_{s}}$, can be found in Cutler et al. (1993); Chernoff and Finn (1993); Zhu et al. (2011).
@@
It can be seen that the SGWB form binary PBHs is below the sensitivity of LIGO O3 but it could be verified by CE and ET in the future. However, the resulting SIGWs exceed the upper limits of NANOGrav 15yr, indicating a discrepancy for PBHs generated by the collapse of Gaussian perturbations to explain SSM200803.
@@ IV Conclusion and Discussion
However, SIGWs are related to PBHs formed by the collapse of overdensities. Given that the merger rate of PBHs is not sensitive with $P(m)$ and it is mainly dependent on $f_{\mathrm{PBH}}$. Therefore, the SGWB from binary PBH mergers would also be consistent with existing observations if the PBHs have other formation mechanisms, such as first-order phase transition Bagui et al. (2023). As a result, the SIGWs would disappear if PBHs had other formation mechanisms. In that case, our results cannot entirely rule out the possibility of PBHs in explaining SSM200803.
@@
Acknowledgments. We thank Prunier et al. for sharing the data on the posterior distribution of SSM200803 Prunier et al. (2023). C.Y. thanks Antonio Riotto, Zu-cheng Chen, Yu-mei Wu and Shi Pi for useful discussions and valuable advice. This work is supported by the National Key Research and Development Program of China Grant No.2020YFC2201502, grants from NSFC (grant No. 11991052, 12047503), Key Research Program of Frontier Sciences, CAS, Grant NO. ZDBS-LY-7009. C.Y. acknowledges the financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855. No 101007855.

### v2

@@ Abstract
In the recent second part of the third observation run by the LIGO-Virgo-KAGRA collaboration, a candidate with sub-solar mass components was reported, which we labelled as SSM200308. This study investigates the premise that primordial black holes (PBHs), arising from Gaussian perturbation collapses, could explain SSM200308. Through Bayesian analysis, we obtain the primordial curvature power spectrum that leads to the merger rate of PBHs aligning with observational data as long as they constitute $f_{\mathrm{PBH}}=5.66^{+58.68}_{-5.44}\times 10^{-2}$ of the dark matter. However, while the gravitational wave (GW) background from binary PBH mergers is within current observational limits, the scalar-induced GWs associated with PBH formation exceed the constraints imposed by pulsar timing arrays, challenging the Gaussian perturbation collapse PBH model as the source of SSM200308.
@@ I Introduction
Since the first detection of a gravitational wave (GW) event by LIGO Abbott et al. (2016a), which marked the dawn of GW astronomy, the LIGO-Virgo-KAGRA collaboration has reported nearly a hundred GW events over the subsequent years.Abbott et al. (2016b, 2018a, 2021, 2024, 2023a). Prior investigations into compact binary coalescences of sub-solar mass (SSM) objects within the collected data yielded no conclusive evidence of such events Abbott et al. (2018b, 2019); Nitz and Wang (2021); Abbott et al. (2022, 2023b); Nitz and Wang (2022); Wang et al. (2024). However, it was not until recently, in the second part of the third observing run Abbott et al. (2023a), that three SSM candidates were reported Abbott et al. (2023b) and they are also analyzed in Morras et al. (2023); Prunier et al. (2023). These three candidates are not claimed as GW events, owing to their relatively large false alarm rates and relatively low signal-to-noise ratio (SNR). Notably, one of the candidates, SSM200308, exhibited a false alarm rate of approximately $0.2$ per year and an $\mathrm{SNR}=8.90$, positioning it on the cusp of claiming a GW detection. With ongoing enhancements in LIGO-Virgo-KAGRA’s sensitivity and the accumulation of observational data, the investigation into the nature and origins of SSM candidates is emerging as a promising and compelling field of study. Recent research has proposed primordial black holes (PBH) as a potential source of SSM candidates Yamamoto et al. (2023); Carr et al. (2024); Crescimbeni et al. (2024), suggesting a novel avenue for understanding these enigmatic phenomena.
PBHs are formed in the very early Universe, originating from the collapse of overmassive regions Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974); Carr (1975). The overabundance of mass in these regions is generated by large curvature perturbations that are enhanced on scales (e.g., Pi et al. (2018); Cai et al. (2018); Cotner and Kusenko (2017); Espinosa et al. (2018); Palma et al. (2020); Pi and Sasaki (2023); Meng et al. (2023)) much smaller than those observed in the cosmic microwave background Aghanim et al. (2020). PBHs are not only promising dark matter (DM) candidates but if a few-thousandth of DM consists of PBHs, they can also explain the GW events observed by the LIGO-Virgo-KAGRA collaboration Sasaki et al. (2016); Chen and Huang (2018); Raidal et al. (2019); De Luca et al. (2020); Hall et al. (2020); Bhagwat et al. (2021); Hütsi et al. (2020); Wong et al. (2021); De Luca et al. (2021); Bavera et al. (2022); Franciolini et al. (2021); Chen et al. (2022); Chen and Hall (2024).
@@ II Bayesian analysis of SSM200308
$\mathcal{R}_{12}(t\mid\!\vec{\theta})$ $=\frac{1.6\times 10^{6}}{\mathrm{Gpc}^{3}\mathrm{yr}}f_{\mathrm{PBH}}^{\frac{53}{37}}\left(\frac{t(z)}{t_{0}}\right)^{-\frac{34}{37}}\eta^{-\frac{34}{37}}\left(\frac{M}{M_{\odot}}\right)^{-\frac{32}{37}}$ (1)
×$\times S\left[M,f_{\mathrm{PBH}},P(m),z\right]P\left(m_{1}\right)P\left(m_{2}\right),$
where $M=m_{1}+m_{2}$, $\eta=m_{1}m_{2}/M^{2}$, $t_{0}$ is the age of the Universe, $f_{\mathrm{PBH}}\equiv\Omega_{\mathrm{PBH}}/\Omega_{\mathrm{CDM}}$ represents the energy fraction of CDM in the form of PBHs and $\sigma_{\mathrm{eq}}\approx 0.005$ Ali-Haïmoud et al. (2017); Chen and Huang (2018) stands for the variance of density perturbations of the rest CDM at matter-radiation equality. $\vec{\theta}$ represents a population of parameters in the PBH mass function. The mass function is normalized as $\int P(m)\mathrm{d}m=1$. The expression of suppression factor, $S$, can be found in Hütsi et al. (2020); Franciolini et al. (2022). We also considered the merger rate contributed by the three-body configurations, which is expressed as Vaskonen and Veermäe (2020); Raidal et al. (2024)
$\mathcal{R}_{12,3}(t\mid\!\vec{\theta})$ $\approx\frac{7.9\times 10^{4}}{\mathrm{Gpc}^{3}\mathrm{yr}}\left(\frac{t}{t_{0}}\right)^{\frac{\gamma}{7}-1}f_{\mathrm{PBH}}^{\frac{144\gamma}{\frac{259}{}}+\frac{47}{37}}$ (2)
$\times$ ${\left[\frac{\langle m\rangle}{M_{\odot}}\right]^{\frac{5\gamma-32}{37}}\left(\frac{M}{2\langle m\rangle}\right)^{\frac{1799}{259}-\frac{2122}{333}}(4\eta)^{-\frac{3\gamma}{7}-1}}$
×$\times\mathcal{K}\frac{e^{-3.2(\gamma-1)}\gamma}{28/9-\gamma}\overline{\mathcal{F}}\left(m_{1},m_{2}\right)P\left(m_{1}\right)P\left(m_{2}\right),$
where $\gamma$ and $\mathcal{K}$ stand for the dimensionless distribution of angular momenta and the hardening of the PBH binary in encounters with other PBHs respectively. We use $\gamma=1$ and $\mathcal{K}=1$ suggested by the numeric simulation Raidal et al. (2019). The $\overline{\mathcal{F}}$ factor is given by
¯$\overline{\mathcal{F}}\left(m_{1},m_{2}\right)$ $\equiv\int_{m\leq m_{1},m_{2}}\mathrm{~{}d}m~{}P(m)\frac{\langle m\rangle}{m}$ (3)
×$\times\left[2\mathcal{F}\left(m_{1},m_{2},m\right)+\mathcal{F}\left(m,m_{1},m_{2}\right)\right],$
with
$\mathcal{F}\left(m_{1},m_{2},m_{3}\right)$ $=m_{1}^{\frac{5}{3}}m_{2}^{\frac{5}{3}}m_{3}^{\frac{7}{9}}\left(\frac{m_{1}+m_{2}}{2}\right)^{\frac{4}{9}}$ (4)
×$\times\left(\frac{m_{1}+m_{2}+m_{3}}{3}\right)^{\frac{2}{9}}\langle m\rangle^{-\frac{43}{9}}.$
@@
where $\mathcal{C}(r)=\mathcal{C}_{g}-\frac{1}{4\Phi}\mathcal{C}_{g}^{2}$ and $\mathcal{C}_{g}=-2\Phi r\zeta^{\prime}(r)$ is linearly related to the comoving curvature perturbation. Here, $\Phi\equiv 3(1+w)/(5+3w)$ and $w$ is the equation of state. In this paper, we consider the softening of the equation of state during the QCD phase transition Musco and Miller (2013); Saikawa and Shirai (2018) where the threshold value of forming PBHs will slightly decrease Borsanyi et al. (2016); Saikawa and Shirai (2018); Franciolini et al. (2022); Musco et al. (2023) since there is less pressure to resist the collapse overdensities. $r_{m}$ in Eq. (5) represents the distance between the location where $\mathcal{C}_{g}$ gains its maximum from and the origin of the peak in curvature perturbation, $\zeta(0)$. The mass of PBH over the horizon mass is given by $m/M_{H}=\kappa\left(\mathcal{C}(r_{m})-\mathcal{C}_{c}\right)^{\gamma}$ Choptuik (1993); Evans and Coleman (1994); Niemeyer and Jedamzik (1998). For simplicity, the scaling parameters $\gamma$ and $\kappa$ during the QCD transition are set to $\kappa=3.3$ and $\gamma=0.36$ Koike et al. (1995). The second-order derivative of the compact function is given by
@@
with $\vec{\lambda}\equiv\{m_{1},m_{2},z\}$, $p(\vec{\lambda}|data)$ is the posterior of SSM200308 obtained in Prunier et al. (2023) and $\beta(\vec{\theta})\equiv\int\mathrm{d}\vec{\lambda}\ VT(\vec{\lambda})\ \mathcal{R}_{12}(\vec{\lambda}|\vec{\theta})$ where $VT(\vec{\lambda})$ is the spacetime sensitivity volume of LIGO-Virgo-KAGRA detectors Abbott et al. (2023b).
@@
We report our analysis in Fig. 1. The median value of the parameters and their $95\%$ equal-tailed credible intervals are $A=0.26^{+0.56}_{-0.13}$, $k_{*}=4.31^{+1.16}_{-0.84}\times 10^{6}\mathrm{Mpc}^{-1}$ and $\sigma=0.42^{+0.44}_{-0.19}$ Moreover, the fraction of PBHs in DM is evaluated to be $f_{\mathrm{PBH}}=5.66^{+58.68}_{-5.44}\times 10^{-2}$ and the local merger rate is given by $\mathcal{R}=3953^{+20939}_{-3371}\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$. The sensitivity volume values from the O3 run in the SSM search for SSM200308 is $\left\langle VT\right\rangle\simeq 3\times 10^{-4}\mathrm{Gpc}^{3}\mathrm{yr}$ Abbott et al. (2023b). Therefore, the expected event numbers are $N=\mathcal{R}\left\langle VT\right\rangle=1.09^{+5.76}_{-0.93}$ which is consistent with Phukon et al. (2021); Abbott et al. (2023b).
Furthermore, we demonstrate the PBH mass function and $f_{\mathrm{PBH}}$ in Fig. 2, together with the median mass of SSM200803. We also show the constraints on $f_{\mathrm{PBH}}$ by EROS/MACHO microlensing at $95\%$ CL Tisserand et al. (2007), OGLE microlensing at $95\%$ CL Mroz et al. (2024a, b), SGWB from binary PBHs by LIGO-Virgo-KAGRA O3 Nitz and Wang (2022) at $90\%$ CL and the LIGO-O3 results in Hütsi et al. (2020) at $2\sigma$ CL by assuming all the O3 events have astrophysical origin, dynamical heating of ultra-faint dwarf galaxies (UFD) Brandt (2016) and accretion constraints of CMB at $95\%$ Ali-Haimoud and Kamionkowski (2017); Aloni et al. (2017); Horowitz (2016); Chen et al. (2016); Poulin et al. (2017). It can be seen from Fig. 2 that the result of current constraints has some overlap with $f_{\mathrm{PBH}}$. As future data accumulate, it may be possible to verify or falsify these SSM PBHs. We remind the readers that all these constraints on $f_{\mathrm{PBH}}$ assume a monochromatic PBH mass function and a precise comparison should generalize these constraints to an extended mass function scenario (see e.g., Carr et al. (2017)). However, the data is in favour of a small width power spectrum, therefore the resulting mass function is approximately monochromatic and the result of $f_{\mathrm{PBH}}$ in Fig. 2 would give readers a rough order of magnitude comparison.
Figure 2: The posterior distribution of the PBH mass function and current constraints on $f_{\mathrm{PBH}}$. The blue bands represent PBH mass function at $95\%$ confidence level. The error bar stands for $f_{\mathrm{PBH}}=5.66^{+58.68}_{-5.44}\times 10^{-2}$
@@ III.2 SGWB form binary PBH coalescences
Here $H_{0}$ is the Hubble value by today, $\Delta=\Omega_{r}(1+z)^{4}+\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}$, $\Omega_{m}$ and $\Omega_{\Lambda}$ are density parameters for matter and dark energy respectively Aghanim et al. (2020). The energy spectrum of a single PBH coalescence event in the source frame, ${dE_{s}}/{df_{s}}$, can be found in Cutler et al. (1993); Chernoff and Finn (1993); Zhu et al. (2011). Note that Eq. (19) also includes the contribution of resolvable events. However, as we will see, the result of SGWB is far below the current upper limits even if it is overestimated.
@@
Although we do not subtract resolvable events, the SGWB still does not conflict with the existing upper limits, it can be seen that the SGWB form binary PBHs is far below the sensitivity of LIGO O3 and it could be verified by CE and ET in the future. However, the resulting SIGWs exceed the upper limits of NANOGrav 15yr, indicating a discrepancy for PBHs generated by the collapse of Gaussian perturbations to explain SSM200803.
@@ IV Conclusion and Discussion
However, SIGWs are related to PBHs formed by the collapse of overdensities. Given that the merger rate of PBHs is not sensitive with $P(m)$ and it is mainly dependent on $f_{\mathrm{PBH}}$. Therefore, the SGWB from binary PBH mergers would also be consistent with existing observations if the PBHs have other formation mechanisms, such as first-order phase transition Bagui et al. (2023). As a result, the SIGWs would disappear if PBHs had other formation mechanisms. In that case, our results cannot entirely rule out the possibility of PBHs in explaining SSM200803. Moreover, the PBH formation model contains uncertainties in calculating $f_{\mathrm{PBH}}$ from the curvature power spectrum. For instance, adopting the peaks theory Bardeen et al. (1986) to evaluate the PBH abundance leads to smaller SIGW Franciolini et al. (2022) and thus alleviates the tension with PTAs.
@@
Acknowledgments. We thank Prunier et al. for sharing the data on the posterior distribution of SSM200803 Prunier et al. (2023). C.Y. thanks Antonio Riotto, Zu-cheng Chen, Yu-mei Wu and Shi Pi for useful discussions and valuable advice. This work is supported by the National Key Research and Development Program of China Grant No.2020YFC2201502, grants from NSFC (grant No. 11991052, 12047503), Key Research Program of Frontier Sciences, CAS, Grant NO. ZDBS-LY-7009.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreements No 101007855 and No 101131233.